(3 P.) The density of graphite is 2.25 g.cm3, and that of diamond is 3.51 g.cm 3. If G0298-2.9 kļ for 1 mol C(graphite) C(diamond), determine the pressure at which graphite and diamond are in equilibrium at 25°C
Answer #1
Free energy = work done
∆G = p∆v
2.9J= P(v2 - v1)
v=m/d
2.9= P(12/2.25 - 12/3.51) × 10^-6(m) × . 10^-3(kg) .

2.9= P(5.33 - 3.418) × 10^-9
2.9= P × 1.91 × 10^-9 
P= 2.9/1.91×10^-9 
P= 1.518 × 10^9 Pa